References

[1] Nicholas Carlini and David Wagner. Towards Evaluating the Robustness of Neural Networks. 2017.

[2] Douglas Crockford and Chip Morningstar. Standard ECMA-404 The JSON Data Interchange Syntax. Dec. 2017. doi: 10.13140/RG.2.2.28181.14560.

[3] Ian J. Goodfellow, Jonathon Shlens, and Christian Szegedy. Explaining and Harnessing Adversarial Examples. 2015.

[4] “IEEE Standard for Robustness Testing and Evaluation of Artificial Intelligence (AI)-based Image Recognition Service”. In: IEEE Std 3129-2023

(2023), pp. 1–34. doi: 10.1109/IEEESTD.2023.10141539.

[5] R.E. Schapire. “The strength of weak learnability”. In: 30th Annual Symposium on Foundations of Computer Science. 1989, pp. 28–33. doi: 10.

1109/SFCS.1989.63451.

[6] Karen Simonyan and Andrew Zisserman. “Very Deep Convolutional Networks for Large-Scale Image Recognition”. In: arXiv 1409.1556 (Sept. 2014).

[7] Christian Szegedy et al. Intriguing properties of neural networks. 2014.

[8] Fisher Yu et al. Deep Layer Aggregation. 2019. eprint: 1707.06484 (cs.CV).